A company currently pays a dividend of $2.8 per share (D0 = $2.8). It is estimated that the company's dividend will grow at a rate of 15% per year for the next 2 years, and then at a constant rate of 7% thereafter. The company's stock has a beta of 1.2, the risk-free rate is 7%, and the market risk premium is 2.5%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Using CAPM,

Required rate of return on stock (Ke) = Risk free rate + Beta x market risk premium

Required rate of return on stock (Ke) = 7% + 1.2 x 2.5%

Required rate of return on stock (Ke) = 10%

Current Dividend (D0) = $2.8

Dividend Growth rate for next 2 years = 15%

Long Term Dividend growth rate (g) = 7%

Dividend for next year that is year 1 (D1) = $2.8 x 115% = $3.22

Dividend for year 2 (D2) = $ 3.22 x 115% = 3.703

Dividend for year 3 (D3) = $ 3.703 x 107% = 3.9622

Therefore value of share at the end of year 2 using Gordon's Dividend Growth Model,

Value of Stock at the end of year 2 = D3 / (ke-g)

Value of Stock at the end of year 2 = 3.9622 / (10%-7%)

Value of Stock at the end of year 2 = $ 132.074

Current Price of stock (P0) = D1/(1+Ke) + D2/(1+Ke)2 + P2/(1+Ke)2

Current Price of stock (P0) = 3.22/(1+0.10) + 3.703/(1+0.10)2 + 132.074/(1+0.10)2

Current Price of stock (P0) = $115.14

Thumbs up please if satisfied. :)

Comment if further doubts in above solution.